Verify maximum Q=f(L,K) subject c=wl + rk
1
Expert's answer
2021-08-25T17:07:47-0400

"Q=f(L,K)"

Q =quantity of output produced.

K= is capital input.

L= is labor input.

F= is a functional form relating the inputs to output. 

"c=wl + rk"

c=wage×labour +interest×capital

rearranging

"K=(\frac{1}{r})c-(\frac{W}{r})L"

"MRTS_{KL}=\frac{W}{r}"